the area of a rectangle is 243cm2. the length is three times greater than the width. what are the dimensions of the rectangle?
Answer by uday1435(57)

Area of rectangle = lw = 3w(w) = 3w^2 = 243 cm^2
w^2 = 243/3 = 81
w = √81 = 9 cms l = 3w = 3(9) = 27
Therefore the length is 27 cms and its width = 9 cms




If you still require further clarifications please write to udayakumar.t.r@gmail.com